OpenStudy (anonymous):

(Help please Medal rewarded and fan) Jenny and her friends are hungry for dessert. They buy $12.25 worth of cookies and ice cream sandwiches. The ice cream sandwiches costs $1.75 each and the cookies cost $0.75 each. If they buy 11 items total, create a system of equations to represent the problem and find the number of cookies and ice cream sandwiches bought by graphing. a) name the variables: x is the number of ice cream sandwiches and y is . b) set up equation for the number of items bought: c) set up equation for the total spent: d) Graph the first equation. Y

OpenStudy (anonymous):

x is sandwiches and y is cookies?

OpenStudy (anonymous):

What is the total number of sandwiches and cookies purchased?

OpenStudy (anonymous):

11 in total

OpenStudy (anonymous):

So, what is the equation for that?

OpenStudy (anonymous):

11= 1.75x+0.75y

OpenStudy (anonymous):

No. In this equation you are only concerned with the number of items purchased.

OpenStudy (anonymous):

So it would be like 5x+6y=11 since the total is 11 and we dont know for sure how many items were bout?

OpenStudy (anonymous):

bought*

OpenStudy (anonymous):

Your doing too much work in your head. x = the number of sandwiches, y = the number of cookies. 5 + 6 = 11, that's true, but not relevant to the question. Sandwiches + cookies = total items purchased.

OpenStudy (texaschic101):

y = cookies x = ice cream sandwiches x + y = 11 (number of items bought ) 1.75x + .75y = 12.25 (total spent) I am not good at graphing

OpenStudy (anonymous):

Your equations look good. rewrite the equations in y = mx + b form to make it easier to graph. In other words, solve for y.

OpenStudy (anonymous):

ok thank you very both of yall it helps

OpenStudy (texaschic101):

sorry that I couldn't help more....gypsy could probably help you with the graphing....she is pretty good at math

OpenStudy (anonymous):

Once you have re-written the equations, I can talk you through a simple process of graphing them.